OpenStudy (geometry_hater):

VERY CONFUSING In the following rectangles, a and b are the sides lengths. Which rectangle has the greatest area? P: a = 12 and b = 3 Q: a = 8 and b = 2 R: a = 7 and b = 6 S: a = 6 and b = 5

OpenStudy (anonymous):

R: its area is a*b=7*6=42 square units

OpenStudy (geometry_hater):

thanks